OpenStudy (ilovecomputers1):

Multiply x to the 3 fourths power times x to the 1 fifth power

OpenStudy (mertsj):

\[x ^{\frac{3}{4}}x ^{\frac{1}{5}}=x ^{\frac{3}{4}+\frac{1}{5}}\]

OpenStudy (ilovecomputers1):

I mean x 3/9

OpenStudy (mertsj):

What is the common denominator for 4 and 5 ?

OpenStudy (ilovecomputers1):

20.

OpenStudy (ilovecomputers1):

Oh I think I see now. So x 15/20 + 4/20 = 19/20. So it would be x 19/20?

OpenStudy (mertsj):

3/4 = ? /20

OpenStudy (mertsj):

yes

OpenStudy (ilovecomputers1):

Thank you so much.

OpenStudy (mertsj):

yw